OpenStudy (anonymous):

Write the expression as a single natural logarithm. 3In x -2 Inc

OpenStudy (raden):

hint: ln(a^b) = b(ln(a))

OpenStudy (anonymous):

ln x^2/c^2

OpenStudy (raden):

no, look the power of x

OpenStudy (anonymous):

idk

OpenStudy (raden):

3In x, should be ln(x^3) isn't it?

OpenStudy (anonymous):

i'm not sure

OpenStudy (raden):

ok, nope.. 3 lnx = ln(x^3) 2 lnc = ln(c^2) subtract both, then use the formula ln(a)-ln(b) = ln(a/b) i sure million percents, u will get this answer.. :D